Beth bought $0.40 stamps and $0.03 stamps at the post office. If she purchased 92 stamps at a total cost of $22, how many of each kind did she buy?
Found 2 solutions by checkley77, rfer:
Answer by checkley77(12844)

.40x+.03(92-x)=22
.40x+2.76-.03x=22
.37x=22-2.76
.37x=19.24
x=19.24/.37
x=52 40 cent stamps were bought.
92-52=40 3 cent stamps were bought.
Proof:
.40*52+.03*40=22
20.8+1.20=22
22=22

Answer by rfer(16322)

x+y=92, x=92-y
.40x+.03y=22
.40(92-y)+.03y=22
36.8-.4y+.03y=22
-.37y=-14.8
y=-14.8/-.37
y=40 $.03 stamps
x=52 $.40 stamps
40*.03=1.20
52*.40=20.80